# Multimodal Prompt Perceiver: Empower Adaptiveness, Generalizability and Fidelity for All-in-One Image Restoration

### 05 Dec 2023 v1

@@ 4.1 Experimental Setup
Settings. Our experiments are conducted in 3 different settings: (1) All-in-one: We train a unified model to solve 10 IR tasks, including deraining, dehazing, desnowing, raindrop removal, low-light enhancement, motion deblurring, defocus deblurring, gaussian denoising, real denoising and challenging mixed degradations removal. (2) Zero-shot: We use the all-in-one pre-trained model to directly solve training-unseen tasks, including under-display camera IR (POLED/TOLED), underwater IR. (3) Few-shot: We fine-tune the all-in-one pre-trained model using a small amount of data (about 3%-5% of the data used by task-specific methods) and adapt it to new tasks, including JPEG compression artifact removal, demosaicking, demoireing.
Datasets and Metrics. For setting(1), a combination of various image degradation datasets is used to evaluate our method, i.e., Rain1400 [22], Outdoor-Rain [50], SSID [35] and LHP [29] for deraining; RESIDE [45], NH-HAZE [5] and Dense-Haze [4] for dehazing; Snow100K [63] and RealSnow [130] for desnowing; RainDrop [74] and RainDS [76] for raindrop removal; LOL-v2 [110] for low-light enhancement; CBSD68, CBSD400 [67], Urban100 [36], Kodak24 [21], McMaster [121] and WED [65] for gaussian denoising; SIDD [1] for real image denoising; GoPro [71] and RealBlur [82] for motion deblurring; and DPDD [2] for defocus deblurring. Considering real-world LQ images may contain more than just a single degradation, we construct a challenging mixed degradation benchmark named MID6, in which the LQ images contain mixed degradations (e.g., low-light&noise&blur, rain&raindrop&noise; See Fig. 7). For setting(2), we utilize TOLED [129] and POLED [129] for under-display camera (UDC) IR. Following [38], 90 pairs from UIEB [47] and 110 pairs from UWCNN [48] are used for underwater IR. For setting(3), we use the first 100 images of DIV2K [3] to fine-tune our method for JPEG compression artifact removal and demosaicking. 5% of the data in TIP2018 [92] training set is used to fine-tune our method for demoireing. We provide a detailed introduction to the datasets used for training and testing in the appendix.
As for evaluation metrics, we adopt PSNR and SSIM as the distortion metrics, LPIPS [122] and FID [32] as the perceptual metrics.
Implementation Details. We use the AdamW optimizer ($\beta_{1}=0.9$, $\beta_{2}=0.999$) with the initial learning rate $1e^{-4}$ gradually reduced to $1e^{-6}$ with the cosine annealing [64] schedule to train our model. The training runs for 500 epochs on 8 NVIDIA A100 GPUs. We adopt DDIM [89] as our sampling strategy (50 steps) for inference. Our model is based on Stable Diffusion 2.1. More details are presented in the appendix.
@@ 4.2 Comparison with state-of-the-art methods
All-in-one. We conduct comparisons between our method and state-of-the-art all-in-one methods as well as task-specific methods. To ensure a fair evaluation, we train all-in-one models from scratch employing our training strategy. Given that Restormer [113] and NAFNet [9] serve as strong general image restoration (IR) baselines, we additionally train both a Restormer model and a NAFNet model within the all-in-one setting.
Table 1 illustrates comprehensive performance comparisons with state-of-the-art (SOTA) methods across 9 tasks. Our method consistently outperforms the compared all-in-one methods on all datasets. Notably, as an all-in-one approach, our method achieves superior results compared to other task-specific methods in the majority of tasks. Additionally, Table 2 presents a comparison on the proposed MID6 benchmark, where our method demonstrates significant advantages in addressing challenging mixed-degraded images. Visual results of some intricate cases from MID6 are presented in Fig. 7, showcasing the enhanced effectiveness of our method in handling mixed degradations. Recognizing the significance of real-world IR challenges, we present more results on real-world datasets in Table 3. Fig. 6 offers visual comparisons across various tasks in real-world scenarios, showcasing the superiority of our method in addressing complex authentic degradations. More comparison results are provided in the appendix.
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Snowy Input / 19.80 dB Restormer [113] / 21.24 dB NAFNet [9] / 20.56 dB AirNet [46] / 21.41 dB TransWeather [94] / 22.26 dB WGWS-Net [130] / 21.29 dB Ours / 23.52 dB GT / PSNR
@@
Figure 6: Real-world visual results on dehazing, desnowing and deraining. Best viewed with zoom in.
@@
Table 9: Ablations of MPerceiver on the Dense-haze [4] dataset.
@@
Baseline (Stable Diffusion) 8.64 0.417 0.802
+Visual Branch 17.21 0.578 0.413
+Visual Branch w/o VP Pool 17.03 0.572 0.416
+(Visual & Textual) Branch 17.44 0.585 0.399
+(Visual & Textual) Branch w/o TP Pool 17.35 0.582 0.405
+(Visual & Textual) Branch + DRM (Full Model) 20.19 0.655 0.394
Zero-shot. As shown in Table 4 and Table 5, we evaluate the performance of each all-in-one method on training-unseen tasks. Our MPerceiver outperforms compared all-in-one methods in all metrics, demonstrating the generalizability of our approach.
Few-shot. As depicted in Table 6, Table 7, and Table 8, we fine-tune the all-in-one methods with limited data to tailor them for new tasks. Notably, our method achieves comparable or superior results compared to task-specific methods trained with substantial amounts of data. This underscores the efficacy of MPerceiver, demonstrating that, post multitask pretraining, it has acquired general representations in low-level vision, allowing for cost-effective adaptation to new tasks.
@@ 4.3 Ablation Study
We perform ablation studies to examine the role of each component in MPerceiver. In Table 9, we initiate with the Stable Diffusion baseline and systematically incorporate or exclude the remaining modules of MPerceiver, including the visual branch, visual prompt (VP) pool, textual branch, textual prompt (TP) pool, and detail refinement module (DRM). The results exhibit a gradual improvement upon the addition of each component and a corresponding decline upon its removal, underscoring the effectiveness of each module. Besides, Fig. 8 visualizes the t-SNE statistics of visual prompt $V^{1}$ and CLIP textual embeddings $\mathcal{E}_{clip}^{txt}(T)$. It demonstrates that our multimodal prompt learning can effectively enable the network to distinguish different degradations. More analysis is provided in the appendix.
@@ 5 Conclusion
This paper introduces MPerceiver, a multimodal prompt learning approach utilizing Stable Diffusion priors for enhanced adaptiveness, generalizability, and fidelity in all-in-one image restoration. The novel dual-branch module, comprising the cross-modal adapter and image restoration adapter, learns holistic and multiscale detail representations. The adaptability of textual and visual prompts is dynamically tuned based on degradation predictions, enabling effective adaptation to diverse unknown degradations. Additionally, a plug-in detail refinement module enhances restoration fidelity through direct encoder-to-decoder information transformation. Across 16 image restoration tasks and 26 benchmarks, including all-in-one, zero-shot, and few-shot scenarios, MPerceiver demonstrates superior adaptiveness, generalizability, and fidelity.

### 20 Mar 2024 v2

@@ 4.1 Experimental Setup
Settings. (1) All-in-one: We train a unified model to solve 10 IR tasks, including deraining, dehazing, desnowing, raindrop removal, low-light enhancement, motion deblurring, defocus deblurring, gaussian denoising, real denoising and challenging mixed degradations removal. (2) Zero-shot: We use the all-in-one pre-trained model to directly solve training-unseen tasks, including under-display camera IR (POLED/TOLED), underwater IR. (3) Few-shot: We fine-tune the all-in-one pre-trained model using a small amount of data (about 3%-5% of the data used by task-specific methods) and adapt it to new tasks, including JPEG compression artifact removal, demosaicking, demoireing.
Datasets and Metrics. For setting(1), a combination of various image degradation datasets is used to evaluate our method, i.e., Rain1400 [24], Outdoor-Rain [53], SSID [38] and LHP [31] for deraining; RESIDE [48], NH-HAZE [6] and Dense-Haze [5] for dehazing; Snow100K [68] and RealSnow [141] for desnowing; RainDrop [84] and RainDS [86] for raindrop removal; LOL-v2 [119] for low-light enhancement; CBSD68, CBSD400 [74], Urban100 [39], Kodak24 [23], McMaster [131], WED [72] and DF2K for gaussian denoising; SIDD [1] for real image denoising; GoPro [80] and RealBlur [92] for motion deblurring; and DPDD [2] for defocus deblurring. Considering real-world LQ images may contain more than just a single degradation, we construct a challenging mixed degradation benchmark named MID6, in which the LQ images contain mixed degradations (e.g., low-light&noise&blur, rain&raindrop&noise; See Fig. 7). For setting(2), we utilize TOLED [140] and POLED [140] for under-display camera (UDC) IR. Following [41], 90 pairs from UIEB [50] and 110 pairs from UWCNN [51] are used for underwater IR. For setting(3), we use the first 100 images of DIV2K [3] to fine-tune our method for JPEG compression artifact removal and demosaicking. 5% of the data in TIP2018 [102] training set is used to fine-tune our method for demoireing. We provide a detailed introduction to the datasets used for training and testing in the Appendix.
We adopt PSNR and SSIM as the distortion metrics, LPIPS [132] and FID [34] as the perceptual metrics, NIQE [78] and BRISQUE [77] as no-reference metrics.
Implementation Details. We use the AdamW optimizer ($\beta_{1}=0.9$, $\beta_{2}=0.999$) with the initial learning rate $1e^{-4}$ gradually reduced to $1e^{-6}$ with the cosine annealing [69] schedule to train our model. The training runs for 500 epochs on 8 NVIDIA A100 GPUs. We adopt DDIM [99] as our sampling strategy (50 steps). Our model is based on SD 2.1. More details are presented in the Appendix.
@@ 4.2 Comparison with state-of-the-art methods
All-in-one. We conduct comparisons between our method and SOTA all-in-one methods as well as task-specific methods. To ensure a fair evaluation, we train all-in-one models from scratch employing our training strategy. Given that Restormer [122] and NAFNet [11] serve as strong general IR baselines, we additionally train both a Restormer and a NAFNet model within the all-in-one setting.
Table 1 illustrates comprehensive performance comparisons with SOTA methods across 9 tasks. Our method consistently outperforms the compared all-in-one methods on all datasets. Notably, as an all-in-one approach, our method even achieves superior results compared to other task-specific methods in many tasks. Additionally, Table 2 presents a comparison on the proposed MID6 benchmark, where our method demonstrates significant advantages in addressing challenging mixed-degraded images. Visual results of some intricate cases from MID6 are presented in Fig. 7, showcasing the enhanced effectiveness of our method in handling mixed degradations. Recognizing the significance of real-world IR challenges, we present more results on real-world datasets in Table 3. Fig. 6 offers visual comparisons across various tasks in real-world scenarios, showcasing the superiority of our method in addressing complex authentic degradations.
@@
Figure 6: Real-world visual results on dehazing and deraining. Best viewed with zoom in.
@@
Table 9: Ablations of MPerceiver. The metrics are reported on the average of deraining, dehazing and raindrop removal.
@@
Baseline (Stable Diffusion) 16.49 0.481 0.538
+Textual Branch 19.19 0.557 0.447
+Textual Branch w/o TP Pool 18.94 0.553 0.457
+Visual Branch 25.05 0.763 0.213
+Visual Branch w/o VP Pool 24.94 0.760 0.216
+(Visual & Textual) Branch 25.31 0.770 0.199
+(Visual & Textual) Branch w/o TP Pool 25.20 0.768 0.206
+(Visual & Textual) Branch w/o VP Pool 25.13 0.767 0.204
+(Visual & Textual) Branch + DRM (Full Model) 29.17 0.842 0.162
Zero-shot. As shown in Tables 4&5, we evaluate the performance of each all-in-one method on training-unseen tasks. MPerceiver outperforms compared methods in all metrics, demonstrating the generalizability of our approach.
Few-shot. As depicted in Tables 6&7&8, we fine-tune the all-in-one methods with limited data to tailor them for new tasks. Notably, our method achieves comparable or superior results compared to task-specific methods trained with substantial amounts of data. This underscores the efficacy of MPerceiver, demonstrating that, post multitask pretraining, it has acquired general representations in low-level vision, allowing for cost-effective adaptation to new tasks.
@@ 4.3 Ablation Study
We perform ablation studies to examine the role of each component in MPerceiver. In Table 9, we initiate with SD and systematically incorporate or exclude the remaining modules of MPerceiver, including the visual branch, visual prompt (VP) pool, textual branch, textual prompt (TP) pool, and detail refinement module (DRM). The results exhibit a gradual improvement upon the addition of each component and a corresponding decline upon its removal, underscoring the effectiveness of each module. Besides, Fig. 8 visualizes the t-SNE statistics of visual prompt $V^{1}$ and CLIP textual embeddings $\mathcal{E}_{clip}^{txt}(T)$. It demonstrates that our multimodal prompt learning can effectively enable the network to distinguish different degradations.
@@ 5 Conclusion
This paper introduces MPerceiver, a multimodal prompt learning approach utilizing Stable Diffusion priors for enhanced adaptiveness, generalizability, and fidelity in all-in-one image restoration. The novel dual-branch module, comprising the cross-modal adapter and image restoration adapter, learns holistic and multiscale detail representations. The adaptability of textual and visual prompts is dynamically tuned based on degradation predictions, enabling effective adaptation to diverse unknown degradations. Additionally, a plug-in detail refinement module enhances restoration fidelity through direct encoder-to-decoder information transformation. Across 16 image restoration tasks, including all-in-one, zero-shot, and few-shot scenarios, MPerceiver demonstrates superior adaptiveness, generalizability, and fidelity.
Acknowledgements: This research is partially funded by Youth Innovation Promotion Association CAS (Grant No. 2022132), Beijing Nova Program (20230484276), National Natural Science Foundation of China (Grant No. U21B2045, U20A20223) and CAAI Huawei MindSpore Open Fund.